Taylor, Moore, and Jenkins are candidates for public office. it is estimated that Moore and Jenkins have about the same probability of winning, and Tayor is believed to be twice as likely to win as either of the others. Find the probability of each candidate winning the election.
Answer by stanbon(75887)

Taylor, Moore, and Jenkins are candidates for public office. it is estimated that Moore and Jenkins have about the same probability of winning, and Tayor is believed to be twice as likely to win as either of the others. Find the probability of each candidate winning the election.
---
Moore prob = x
Jenkins prob = x
Taylor probl = 2x
------
Equation:
x + x + 2x = 1
4x = 1
x = 1/4 (Prob of Moore and of Jenkins)
2x = 1/2 (Prob of Taylor)
=======================================
Cheers,
Stan H.